# MedS3: Towards Medical Small Language Models with Self-Evolved Slow Thinking

### v1

@@ 2.3 PRM Fine-tuning
A satisfactory PRM should have the following properties: (1) For a set of reasoning steps starting from the same reasoning trace, it should differentiate between each step through their rollout values; (2) for a correct leaf and incorrect leaf, it should tend to assign a large value for the correct leaf and a low value for the incorrect leaf.
@@
Therefore, we gather all finish nodes and their corresponding traces: $\{(T_{k}^{j},[s_{0},s_{1},\cdots,s_{k}^{j}],v_{k}^{j})\mid k=1,2,\cdots,K;j=1,2\cdots,J\}$, where $K$ is the max depth of the evolved tree, $J$ is the max width of the $j$-th layer, and $v_{k}^{j}\in(0,1)$. The reasoning traces $[s_{0},s_{1},\cdots,s_{k}^{j}]$ are assembled with “Step k:”. This naturally enables $V_{\theta}$ to distinguish between siblings inheriting from the same parent node: $T_{k}^{j_{1}}$ and $T_{k}^{j_{2}}$ with distinct rollout values $v_{k}^{j_{1}}$ and $v_{k}^{j_{2}}$. To realize the second property, $V_{\theta}$ should not be biased by the distribution of value 0 and value 1, since a big difference between these two sparse values tends to mislead the $V_{\theta}$ to assign either 0 or 1 for all leaf nodes, instead of learning the intrinsic semantic correctness. To this end, we tallied the correct and incorrect finish nodes: $\{T_{k}^{0}\mid v_{T_{k}}=0\}$ and $\{T_{k}^{1}\mid v_{T_{k}}=1\}$, and randomly sampled elements $\min\left(|\{T_{k}^{0}\}|,|\{T_{k}^{1}\}|\right)$ from these two sets, maintaining a balanced distribution of the nodes of correct and incorrect reasoning: $\{\hat{T}_{k}^{0}\}$ and $\{\hat{T}_{k}^{1}\}$. The sampled nodes are combined and form the final $V_{\theta}$ tuning set: $D_{V_{\theta}}=\{\hat{T}_{k}^{0}\}\cup\{\hat{T}_{k}^{1}\}$.
@@
Previous works in the Math domain choose to directly learn the rollout value (Zhang et al., 2024a) or learn the pair-wise ranking preference (Guan et al., 2025). However, in our work, we choose to learn the prediction of the correctness probability of an intermediate step using a 2-class cross-entropy loss. The PRM $V_{\theta}$ is tuned using the same policy model architecture, with the language model head replaced by a token classification layer with a cross-entropy loss at the end of each step. Although Zhang et al. (2025) suggests that the PRM label should be set to 1 (a hard label) once the rollout score is above zero, empirically we borrow the insight from label-smoothing (Szegedy et al., 2016), where a soft label groups positive internal steps tightly in the representation space. Specifically, we set the label $y_{k}$ as the original node rollout value, which is a float number between 0 and 1, and optimize the $V_{\theta}$ using the following loss function:
@@
where $\hat{y}_{k}$ is the predicted probability of the given step and $y_{k}$ is the label. This soft-label training, not only encourages $V_{\theta}$ to cluster preferred and dis-preferred steps, but also prevents the learning of fuzzy labels (rollout value around 0.5)
@@ 2.4 PRM-guided Decoding
A well-trained PRM can serve as an oracle quality verifier, to select both suitable intermediate reasoning steps as well as complete reasoning trajectories. Referring to the inference scaling law (Wu et al., 2024d), enlarging the inference token budget is a deterministic way to enhance the downstream reasoning performance. To balance the advantages of enlarging the inference budget and unwillingly reduced efficiency, we propose a new decoding strategy that relies on the well-configured PRM to select the output from sampled candidates.
@@
[…]
@@
where $n\leq M$ is the number of answers with different semantics obtained from the $M$ responses, $v_{a_{n}}=\sum_{m}\mathds{1}_{a_{y_{m}}=a_{n}}\cdot v_{y_{m}}$ is the sum of values of all semantically equal traces.
@@ 3 Data Statistics
A slow-thinking system in medical scenarios should both excel at answering exam questions, but also handling real-world clinical scenarios. To this end, we extend beyond traditional multiple-choice questions, and incorporate more diverse datasets into the seed evolution datasets. We gather four types of clinical tasks, with each task containing different medical datasets:
  1. 1.
    Long Context Question Answering: This dimension is to enable MedS3 to capture useful information from the given context and use long-chain reasoning to answer problems. This dimension covers BioMRC (Pappas et al., 2020), HeadQA Topic Classification (Vilares and Gómez-Rodríguez, 2019; Wu et al., 2024b), and DDX-Plus (Tchango et al., 2022).
  2. 2.
    Knowledge-Intensive Question Answering: This dimension is expected to enable MedS3 to use long-chain reasoning to answer knowledge-intensive problems. This dimension covers MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and PubMedQA (Jin et al., 2019).
  3. 3.
    General Question Answering in Medical: This part leverages general data in bio-medicine domains to enhance the generality of MedS3, which includes SciQ (Welbl et al., 2017), Evidence Inference (DeYoung et al., 2020) and Head QA (Vilares and Gómez-Rodríguez, 2019).
  4. 4.
    Medical Natural Language Inference: This dimension is considered to opt MedS3 to discriminate biomedical research concepts and corresponding descriptions. This dimension covers Healthfact Classification (Kotonya and Toni, 2020), Medical Question Pair (MQP; McCreery et al. (2020)), and catalonia-independence-corpus (CIC; Zotova et al. (2020)).
  5. 5.
    Diagnosis Question Answering: This dimension is related to real-world clinical scenarios, including disease diagnosis and classification and drug related questions. We choose Covid-19 Classification (Lab, 2020), Drug-Dose Extraction and Adverse Drug Event Classification (Huynh et al., 2016; Wu et al., 2024b).
For each involved dataset, we randomly select 500 items and form a seed dataset with 8,000 instances. We show the dataset statistics in Fig. 2. For more detailed description of each dataset, please refer to Appendix B.
Knowledge QA General Medical QA Long-Context QA Diagnosis QA NLI
Models MedQA MedMCQA HealthFact PubMedQA BioASQ MMLU BioMRC DDXPlus PubHealth DrugDose SEER Avg.
@@
Qwen2.5-7b 55.54 54.12 52.69 53.40 73.62 74.38 56.48 31.25 57.11 60.87 33.07 54.78
Llama3-8b 57.50 55.92 70.88 56.40 75.73 68.55 56.50 35.30 64.09 73.91 47.07 60.17
Llama3.1-8b 61.51 57.42 63.97 59.00 71.36 72.52 55.60 19.00 61.82 73.91 52.62 58.98
QwQ 68.89 61.03 66.08 48.60 73.62 74.18 79.76 45.40 63.36 39.13 37.26 59.76
@@
MMedS-Ins 53.57 48.24 69.64 56.60 77.35 50.86 31.47 97.53 54.26 95.65 97.93 66.65
MedLlama3 55.85 59.36 68.10 66.40 84.63 70.08 47.97 22.50 62.39 69.57 50.69 59.78
Med42 50.20 49.70 81.57 55.40 74.76 61.43 57.26 31.35 59.14 65.22 37.14 56.65
OpenBioLLM 50.20 50.56 53.28 41.40 47.73 61.69 27.46 16.55 18.77 34.78 46.48 40.81
UltraMedical3-8b 68.89 61.82 72.73 51.60 80.58 75.08 45.18 36.70 66.13 60.87 24.55 58.56
UltraMedical3.1-8b 70.93 62.78 70.20 56.40 77.18 76.43 54.26 31.55 59.14 56.52 45.86 60.11
@@
HuatuoGPT-o1 62.53 59.31 66.08 69.20 87.70 70.53 50.98 40.20 24.61 56.52 46.85 57.68
MedS3
    COT 65.91 60.55 64.73 56.80 78.48 75.66 55.84 51.65 57.03 73.91 48.97 62.68
    SC 70.93 64.21 70.37 58.20 79.13 79.63 63.66 57.00 64.42 86.96 52.19 67.88
    PRM 71.88 65.20 79.97 59.60 80.10 79.50 77.12 65.20 73.03 95.65 58.36 73.24
Table 2: Experiment results in 11 medical datasets among four types of models. We highlight the best results with bold and underlines the second-best results. MedS3 achieves superior performances on real-world clinical datasets, especially in DrugDose and DDX-Plus.
@@ 4.1 Experiment Setups
We choose Llama3.1-8B-Instruct as the initialization of both the policy model and PRM model. We select MedQA (Jin et al., 2021), PubMedQA (Jin et al., 2019), MedMCQA (Pal et al., 2022), PubHealth (Kotonya and Toni, 2020), BioMRC (Pappas et al., 2020), HealFact Classification (Kotonya and Toni, 2020), Drug Dose Extraction (Huynh et al., 2016), DDX-Plus (Tchango et al., 2022) as the in-domain evaluation benchmarks, the medical parts of MMLU (Hendrycks et al., 2021), BioASQ (Tsatsaronis et al., 2012) and SEER Classification (Dubey et al., 2023) as the out-of-domain evaluation sets. We provide three different decoding strategies for MedS3, including CoT, Self-Consistency and our proposed PRM guided Vote-Sum. COT (Wei et al., 2022) directly prompts models to generate a long reasoning chain and outputs the answer with “The answer is {answer}” for the convenience of answer extraction. Self-Consistency (Wang et al., 2023) generates $N=16$ samples for a given problem, and we select the one whose answer appears most times among the $N$ outputs. For PRM-guided Vote-Sum, we generate $N=12$ samples for datasets where SC has already achieved a high score (MedMCQA, Med-MMLU and BioASQ) and generate $N=16$ samples for other datasets. This is reasonable since not all generated samples are of high quality due to the sampling parameters top_p. We manually filter some low-probability samples to prevent them from disturbing the Vote-Sum computation. The hyperparameters of synthesis, self-training and evaluation are presented in Appendix C.
@@
We choose the following categories to serve as baselines: (1) Proprietary general LLMs, including GPT-3.5-turbo (OpenAI, 2022) and GPT-4o-mini (OpenAI, 2023); (2) Open-sourced general LLMs, including Llama 3 8B, Llama 3.1 8B (Dubey et al., 2024) and Qwen2.5 7B (Yang et al., 2024), and QWQ-preview-32B (Team, 2024); (3) Open-sourced Medical LLMs, including MedLlama 3 8B [^2] https://huggingface.co/ProbeMedicalYonseiMAILab/medllama3-v20, MMedS-Ins-Llama-3-8B (Wu et al., 2024b), Med42 (Christophe et al., 2024), OpenBioLLM (Ankit Pal, 2024), and UltraMedical3-8B and UltraMmedical3.1-8B (Zhang et al., 2024b). We also compare our method with HuatuoGPT-o1-8B (Chen et al., 2024).
@@ 4.2 Main Results
We present the comprehensive experiment results in Table 2. The results unveil that most prior medical LLMs show superior results in traditional medical benchmarks (MedQA or PubMedQA); while such superiority cannot generalize to out-of-distribution real-world clinical benchmarks (DDXPlus or DrugDose), which results in their sub-optimal overall performance compared to Llama3-8B. This suggests that without choices as prior knowledge, current medical models have a limited application scope. In contrast, our MedS3 is not optimized exclusively for multiple-choice medical datasets; therefore, MedS3 achieves the best performance among all open-sourced competitions. As an 8B model, MedS3 achieves +13.07 average performance gains with respect to the base model in the overall assessment, which not only outperforms medical-oriented models as well as GPT-4o-mini with the PRM as a supervisor. Although MedS3 lags behind on BioASQ and PubmedQA compared to HuatuoGPT and MedLlama3, MedS3 shows superior performance on reasoning-intensive benchmarks, including MedQA and MedMCQA, as well as clinical benchmarks. This verifies that MedS3 learns medical reasoning philosophies and the clinical deduction process. Another model MMedS-Ins, which focused on clinical usage, directly post-pretraining and fine-tuning on millions of medical corpus, harvests superior performance on in-domain clinical tasks, such as SEER and DDXPlus. However, directly fine-tuning on question-answer pairs inevitably makes the model loss the ability to output long responses, which is extremely important in other reasoning tasks. As a result, although the over-fitting model achieves superior performance in three clinical benchmarks, our model MedS3, fine-tuned with a self-evolve pipeline, hardly impairs the internal distribution and hence possesses comprehensive clinical usage.
@@ 5 Analysis
Settings Traditional QA Clinical Tasks Mean
MedQA MedMCQA Healthfact DDXPlus SEER
Base 61.51 57.42 63.97 19.00 52.62 50.90
BoN-min 67.64 59.20 73.57 52.40 41.24 58.81
BoN-prod 65.51 60.00 75.51 57.30 47.67 61.20
BoN-last 60.17 58.00 78.37 65.10 61.67 64.66
PRM-Vote-Sum 71.01 65.20 79.97 65.20 58.3 67.95
Table 3: Comparison of Vote-Sum with other methods utilizing PRM under the same token budgets. Our proposed vote-sum achieves the best performance, while recently proposed BoN-last is secondary to us.
5.1 Selection of Vote-Sum
In this section, we compare our proposed decoding method PRM-guided Vote sum to the previously widely adopted PRM decoding method: Best-of-N (BoN). We compare three variants of BoN, including BoN-min (Lightman et al., 2023), BoN-prod (Lightman et al., 2023) and BoN-last (Zhang et al., 2025) The first method, BoN-min, takes the minimum step value of a reasoning trajectory as the estimation of the whole sequence. The second method, BoN-prod, takes the production of step values of a whole trajectory as the sequence value prediction. The last method is recently proposed, which deems that PRM trained with Monte-Carlo-rollout (MC-rollout) values are basically an ORM and hence only the last step’s prediction is authentic. We select two traditional QA tasks and three clinical tasks to conduct comparison and show results in Table 3. A very interesting finding is that in two datasets of traditional QA, BoN-last achieves extremely bad results, but shows great performance on par with Vote-Sum in clinical tasks. We hypothesize that clinical tasks and Olympia-level math problems are both difficult for Llama3.1-8B, and hence the findings that PRM behaves more like an ORM transfer successfully to clinical tasks. On the other hand, on relatively advantageous tasks, MC-rollout value encodes values of both future and current steps, which leads the PRM to perform as a human-labeled PRM to distinguish a bad step from candidates. As a result, in traditional medical tasks, we use the minimum value of a whole sequence to represent the whole sequence’s correctness, while in clinical tasks, we use the last step’s value to serve as the whole response’s score.
Figure 3: Overview of the used seed datasets.

### v2

@@ 2.3 PRM Fine-tuning
We gather all finish nodes and their corresponding traces: $\{(T_{k}^{j},[s_{0},s_{1},\cdots,s_{k}^{j}],v_{k}^{j})\mid k=1,2,\cdots,K;j=1,2\cdots,J\}$, where $K$ is the max depth of the evolved tree, $J$ is the max width of the $j$-th layer, and $v_{k}^{j}\in(0,1)$. The reasoning traces $[s_{0},s_{1},\cdots,s_{k}^{j}]$ are assembled with $t_{s}$. This naturally enables $V_{\theta}$ to distinguish between siblings inheriting from the same parent node: $T_{k}^{j_{1}}$ and $T_{k}^{j_{2}}$ with distinct rollout values $v_{k}^{j_{1}}$ and $v_{k}^{j_{2}}$. To prevent $V_{\theta}$ from being biased by the distribution of value 0 and value 1, we tallied the correct and incorrect finish nodes: $\{T_{k}^{0}\mid v_{T_{k}}=0\}$ and $\{T_{k}^{1}\mid v_{T_{k}}=1\}$, and randomly sampled elements $\min\left(|\{T_{k}^{0}\}|,|\{T_{k}^{1}\}|\right)$ from these two sets, maintaining a balanced distribution of the nodes of correct and incorrect reasoning: $\{\hat{T}_{k}^{0}\}$ and $\{\hat{T}_{k}^{1}\}$. The sampled nodes are combined and form the final $V_{\theta}$ tuning set: $D_{V_{\theta}}=\{\hat{T}_{k}^{0}\}\cup\{\hat{T}_{k}^{1}\}$.
@@
Previous works in the math domain choose to directly learn the rollout value (Zhang et al., 2024a) or learn the pair-wise ranking preference (Guan et al., 2025). However, in our work, we choose to learn the prediction of the correctness probability of an intermediate step using a 2-class cross-entropy loss. The PRM $V_{\theta}$ is tuned based on the tuned policy model, with the language model head replaced by a token classification layer with a cross-entropy loss at the end of each step. Although Zhang et al. (2025) suggests that the PRM label should be set to 1 (a hard label) once the rollout score is above zero, empirically we borrow the insight from label-smoothing (Szegedy et al., 2016), where a soft label groups positive internal steps tightly in the representation space. We show the comparison between the soft and hard label in Appendix B.1. Specifically, we set the label $y_{k}$ as the original node rollout value, and optimize $V_{\theta}$ using the following loss function:
@@
where $\hat{y}_{k}$ is the predicted probability of the given step and $y_{k}$ is the label. This soft-label training, not only encourages $V_{\theta}$ to cluster preferred and dis-preferred steps, but also prevents the learning of fuzzy labels (rollout value around 0.5).
Figure 2: Overview of the PRM-guided Vote-Sum.
@@ 2.4 PRM-guided Decoding
Referring to the inference scaling law (Wu et al., 2024d), enlarging the inference token budget is a deterministic way to enhance the downstream reasoning performance. To balance between the advantages of enlarging the inference budget and unwillingly reduced efficiency, we propose a new decoding strategy named PRM guided Vote-Sum that considers both response estimation and semantic coherence, which is illustrated in Fig. 2.
@@
This method employs $\pi$ to sample $M$ candidate responses $\{y_{m}\mid m=1,2,\cdots,M\}$ given an input problem $x$ and uses the PRM $V_{\theta}$ to select the response whose answer is estimated to have the highest values in total. Specifically, for a simple output $y_{m}$, we split it into $K$ steps $\{s_{m}^{k}\mid k=1,2,\cdots,K\}$ with pre-defined coherence phrase $t_{s}$. The PRM assigns a score $q\in[0,1]$ for each internal step $\{q_{m}^{k}=V_{\theta}(s_{m}^{1}\oplus s_{m}^{2}\oplus\cdots\oplus s_{m}^{k})\mid k=1,2,\cdots,K\}$. The overall response score $v_{y_{m}}$ takes the minimum value of the score chain (Lightman et al., 2023) or the last value (Zhang et al., 2025). The Vote-Sum strategy comprehensively considers the occurrence of semantically equivalent outputs but also the confidence score predicted by $V_{\theta}$:
@@
where $n\leq M$ is the number of different answers obtained from the $M$ responses, $v_{a_{n}}=\sum_{m}\mathds{1}_{a_{y_{m}}=a_{n}}\cdot v_{y_{m}}$ is the sum of values of samples whose answers equal to $a_{n}$.
@@ 3 Data Statistics
A slow-thinking system in medical scenarios should both excel at exam-level question answering (QA) and handling real-world clinical scenarios, like diagnosis (Tchango et al., 2022), specific disease syndrome (Lab, 2020) and drug-related problems (Huynh et al., 2016). However, previous works only focused on a simple scenario, with only limited data diversity, especially multiple-choice QA, to train reasoning models. To approximate realistic clinical usage and promote medical reasoning models on a diverse range of clinical tasks, we curate a training corpus, from 16 existing public medical datasets and divide them into five dimensions according to the task category. For each dataset, we randomly select 500 items and form a seed dataset with around 8,000 instances. We show the dataset statistics in Fig. 3. The details about the definition of the five dimensions and the corresponding tasks can be found in Appendix E.
Models Knowledge QA BioMedical QA Long Context QA NLI Diagnosis QA Avg.
MedQA MedMCQA PubMedQA BioASQ MMLU BioMRC DDXPlus PubHealth HealthFact DrugDose SEER
@@
Qwen2.5-7B 55.54 54.12 53.40 73.62 74.38 56.48 31.25 57.11 52.69 60.87 33.07 54.78
Llama3-8B 57.50 55.92 56.40 75.73 68.55 56.50 35.30 64.09 70.88 73.91 47.07 60.17
Llama3.1-8B 61.51 57.42 59.00 71.36 72.52 55.60 19.00 61.82 63.97 73.91 52.62 58.98
R1-Distill-Llama8B 50.12 48.89 46.60 70.55 68.42 53.49 36.10 55.73 62.04 69.57 31.71 53.93
QwQ-32B-preview 68.89 61.03 48.60 73.62 74.18 79.76 45.40 63.36 66.08 39.13 37.26 59.76
R1-Distill-Qwen32B 76.83 66.27 38.20 78.32 85.07 78.66 53.90 59.95 63.80 82.61 26.22 64.53
@@
MMedS-Ins 53.57 48.24 56.60 77.35 50.86 31.47 97.53 54.26 69.64 95.65 97.93 66.65
MedLlama3 55.85 59.36 66.40 84.63 70.08 47.97 22.50 62.39 68.10 69.57 50.69 59.78
Med42 50.20 49.70 55.40 74.76 61.43 57.26 31.35 59.14 81.57 65.22 37.14 56.65
OpenBioLLM 50.20 50.56 41.40 47.73 61.69 27.46 16.55 18.77 53.28 34.78 46.48 40.81
UltraMedical3-8B 68.89 61.82 51.60 80.58 75.08 45.18 36.70 66.13 72.73 60.87 24.55 58.56
UltraMedical3.1-8B 70.93 62.78 56.40 77.18 76.43 54.26 31.55 59.14 70.20 56.52 45.86 60.11
@@
HuatuoGPT-o1 62.53 59.31 69.20 87.70 70.53 50.98 40.20 24.61 66.08 56.52 46.85 57.68
MedS3 (Ours)
    CoT 65.91 60.55 56.80 78.48 75.66 55.84 51.65 57.03 64.73 73.91 48.97 62.68
    SC 70.93 64.21 58.20 79.13 79.63 63.66 57.00 64.42 70.37 86.96 52.19 67.88
    P-VS 71.88 65.20 59.60 80.10 79.50 77.12 65.20 73.03 79.97 95.65 58.36 73.24
Table 2: Experiment results in 11 medical datasets among four types of models. We highlight the best results with bold and underlines the second-best results. MedS3 with PRM guided Vote Sum (P-VS) achieves superior performances on real-world clinical datasets.
@@ 4.1 Experiment Setups
We choose Llama3.1-8B-Instruct as the initialization of both the policy model and PRM model. We select MedQA (Jin et al., 2021), PubMedQA (Jin et al., 2019), MedMCQA (Pal et al., 2022), PubHealth (Kotonya and Toni, 2020), BioMRC (Pappas et al., 2020), HealFact Classification (Kotonya and Toni, 2020), Drug Dose Extraction (Huynh et al., 2016), DDX-Plus (Tchango et al., 2022) as the in-domain evaluation benchmarks, the medical parts of MMLU (Hendrycks et al., 2021), BioASQ (Tsatsaronis et al., 2012) and SEER Classification (Dubey et al., 2023) as the out-of-domain evaluation sets. We provide three different decoding strategies for MedS3, including CoT (Wei et al., 2022), Self-Consistency (SC) and our proposed P-VS. The hyperparameters of synthesis, self-training and evaluation are presented in Appendix F.
@@
We choose the following categories to serve as baselines: (1) Proprietary general LLMs, including GPT-3.5-turbo (OpenAI, 2022) and GPT-4o-mini (OpenAI, 2023); (2) Open-sourced general LLMs, including Llama 3 8B, Llama 3.1 8B (Dubey et al., 2024) and Qwen2.5 7B (Yang et al., 2024), QWQ-preview-32B (Team, 2024) and two distilled models from DeepSeek R1 (Guo et al., 2025): R1-Distill-Llama8B and R1-Distill-Qwen32B; (3) Open-sourced Medical LLMs, including MedLlama 3 8B [^2] https://huggingface.co/ProbeMedicalYonseiMAILab/medllama3-v20, MMedS-Ins-Llama-3-8B (Wu et al., 2024b), Med42 (Christophe et al., 2024), OpenBioLLM (Ankit Pal, 2024), and UltraMedical3-8B and UltraMedical3.1-8B (Zhang et al., 2024b). We also compare our method with HuatuoGPT-o1-8B (Chen et al., 2024).
@@ 4.2 Main Results
We present the comprehensive experiment results in Table 2. The results unveil that most prior medical LLMs show superior results in traditional medical benchmarks (MedQA or PubMedQA); while such superiority cannot generalize to out-of-distribution real-world clinical benchmarks (DDXPlus or DrugDose), which results in their sub-optimal overall performance compared to Llama3-8B. In contrast, our MedS3 is not optimized exclusively for multiple-choice medical datasets and hence achieves the best performance among all open-sourced competitions. As an 8B system, MedS3 achieves +13.07 average performance gains with respect to the base model in the overall assessment, which not only outperforms medical-oriented models as well as general reasoning models.
Specifically, compared to HuatuoGPT-o1 and MedLlama3, MedS3 shows superior performance on reasoning-intensive benchmarks, including MedQA and MedMCQA, as well as clinical benchmarks. This verifies that MedS3 learns medical reasoning philosophies and the clinical deduction process. Another model MMedS-Ins, which directly post-pretrains and fine-tunes on millions of clinical corpus, harvests superior performance on in-domain clinical tasks, such as SEER and DDXPlus. However, directly fine-tuning on question-answer pairs inevitably makes the model lose the ability to output long responses, which is extremely important in reasoning tasks. In contrast, our model MedS3, possesses a comprehensive performance on both traditional tasks and clinical scenarios with strong reasoning abilities.
@@ 5 Analysis
Select Model Value Estimation Traditional QA Clinical Tasks
Majority ORM PRM Min Prod Last Vote-Sum MedQA MedMCQA Healthfact DDXPlus SEER Mean
70.93 64.21 70.37 57.00 52.19 62.94
67.87 60.63 63.38 53.75 48.71 58.87
71.88 64.67 70.12 57.15 53.44 63.45
67.64 59.20 73.57 52.40 41.24 58.81
65.51 60.00 75.51 57.30 47.67 61.20
60.17 58.00 78.37 65.10 61.67 64.66
71.88 65.20 79.97 65.20 58.36 68.12
Table 3: Comparison of P-VS with other decoding methods under the same token budgets. “Majority” means we use SC to select the final response. “ORM with Last” means selecting the response with the max ORM value.
5.1 Effectiveness of P-VS
In this section, we compare our proposed decoding method P-VS to the previously widely adopted PRM decoding method: Best-of-N (BoN) as well as ORM guided methods (ORM with Last and ORM with Vote-Sum). We compare three variants of BoN, including BoN-min (Lightman et al., 2023) which takes the minimum step value of a reasoning trajectory as the estimation of the whole sequence, BoN-prod (Lightman et al., 2023) which takes the production of step values instead, and BoN-last (Zhang et al., 2025), which treats the PRM as an ORM. We select two traditional QA tasks and three clinical tasks to conduct comparison and show results in Table 3. A very interesting finding is that in the two traditional QA datasets, BoN-last achieves extremely bad results, but shows great performance on par with Vote-Sum in clinical tasks. We hypothesize that clinical tasks and Olympia-level math problems are both difficult for Llama3.1-8B, and hence the findings that PRM behaves more like an ORM transfer successfully to clinical tasks. On the other hand, on relatively advantageous tasks, MC-rollout value encodes values of both future and current steps, which leads the PRM to perform as a human-labeled PRM to distinguish steps precisely. As a result, in traditional medical tasks, we use the minimum value among step values to represent the whole sequence’s correctness, while in clinical tasks, we use the last step’s value instead. Compared to ORM-based methods, they only excel at traditional QA benchmarks, lagging far behind all PRM-based methods in clinical usage, further validating the efficacy of process supervision. ORM with Vote-Sum surpasses ORM with last by a large margin, which also verifies the effectiveness of the proposed Vote-Sum strategy. We present the error rate of PRM in Appendix B.2 to show that rule-verified PRM is sufficient in detecting erroneous clinical reasoning steps.
Figure 4: Budget scaling on clinical benchmarks. When the token budget increases, MedS3’s improves rapidly and then slows down as the budget is further increased.